1) Let L1 and L2 be lines defined by :

x = w0 + su, s e R(element of R)
y= w1 + tv, t e R(element of R) respectively

Show that L1 and L2 are parallel if and only if u = kv for some k e R. (element of R)

2) Find the plane that passes through the point (2,4,-3) and is parallel to the plane - 2x + 4y - 5z + 6 = 0

3) Find the line that passes through the point (2,5,3) and is perpendicular to the plane 2x - 3y + 4z + 7 = 0

4) Find an equation of the plane passing through the point (-2,3,4) and is perpendicular to the line passing through the points (4,-2,5) and (0,2,4).
1
Expert's answer
2020-04-27T19:24:42-0400

1.) Let L1 and L2 be lines defined by :

"x = w\omicron + su, s\in \real"

"y= w1 + tv, t\in R\phantom{i} respectively"

Show that L1 and L2 are parallel if and only if u = kv for some "k \in R"


Let L1 and L2 be lines defined by

"x = w\omicron + su, s\in \real" ....................................1

and

"y= w1 + tv, t\in R" ,..................................2

respectively.

Show that L1 and L2 are parallel if and only if "u = kv" for some "k\in R" .

PROPOSITION

L1 AND L2 ARE PARALLEL

we show that:

"u=kv"

for L1 and L2 to be parallel .

their directional vector should be scalar multiple of each other.

directional vector of L1 is given by "su"

directional vector of L2 is given by "tv"

then if L1 and L2 are parallel , "su" and "tv" shall be scalar multiple of each other

that is:

"su=m*tv" , where m is scalar

substituting "k=m*t\/s" to the equation "u=kv" then k is a scalar since t,s and m are all scalar,

therefore, L1 and L2 are parallel since "u=kv" and k is the scalar multiple, u and v are directional vectors of L1 and L2.

2.) Find the plane that passes through the point "(2,4,\u22123)" and is parallel to the plane "- 2x + 4y - 5z + 6 = 0"

the equation of the plane parallel to "- 2x + 4y - 5z + 6 = 0" is given be

"- 2x + 4y - 5z + c = 0"

since it passes through the point "(2,4,-3)" . to obtain c we substitute the point to the equation

"- 2x + 4y - 5z + c = 0\\
-2(2)+4(4)-5(-3)+c=0\\-4+16+15+c=0\\\implies c=-27"

hence the equation of the plane parallel to "- 2x + 4y - 5z + 6 = 0" is


"- 2x + 4y - 5z -27 = 0"


3.) Find the line that passes through the point"(2,5,3)" and is perpendicular to the plane "2x - 3y + 4z + 7 = 0"

the equation perpendicular to "2x - 3y + 4z + 7 = 0" is given by

"\dfrac{x-x\omicron}{2}=\dfrac{y-y\omicron}{-3}=\dfrac{z-z\omicron}{4}"

and since the line has pass through "(2,5,3)" then the equation of the line perpendicular to "2x - 3y + 4z + 7 = 0" will be

"\dfrac{x-2}{2}=\dfrac{y-5}{-3}=\dfrac{z-3}{4}"


4.) Find an equation of the plane passing through the point (-2,3,4) and is perpendicular to the line passing through the points (4,-2,5) and (0,2,4).

  • solution

The line passing through "(4,-2,5)" and "(0,2,4)" has the directional ratio of

"<4-0,-2-2,5-4>\\=<4,-4,1>"

hence the equation of the plane passing through "(x\omicron,y\omicron,z\omicron)" and the directional ratio of the line perpendicular to the plane

"<a,b,c>"

is given by

"a(x-x\omicron)+b(y-y\omicron)+c(z-z\omicron)=0"

then:

the required equation of the plane passing through "(-2,3,4)" and the directional ratio of the line perpendicular to the plane "<4,-4,1>" is given by:


"4(x+2)-4(y-3)+1(z-4)=0\\\implies 4x+8-4y+12+z-4=0\\putting\phantom{i}like\phantom{i}terms\phantom{i}together:""4x-4y+z+16=0"